A researcher wanted to determine whether a program for routine cholesterol screening for college men
would identify high serum cholesterol levels. The researcher decided that above 200 would be high. Twenty
students were randomly selected for cholesterol screening, and the results were given below. Is there any
evidence that asymptomatic college men have elevated serum cholesterol values? Would a large-scale screening program be warranted? Use a = 0.05
160 210 200 230 210 150 220 190 210 220
160 220 240 180 200 200 210 170 180 220

Answer by stanbon(75887)

A researcher wanted to determine whether a program for routine cholesterol screening for college men would identify high serum cholesterol levels.
The researcher decided that above 200 would be high.
Twenty students were randomly selected for cholesterol screening, and the results were given below.
Is there any evidence that asymptomatic college men have elevated serum cholesterol values?
Would a large-scale screening program be warranted? Use a = 0.05
160 210 200 230 210 150 220 190 210 220
160 220 240 180 200 200 210 170 180 220
----
x-bar = 199
s = 25.11
z(199) = (200-199)/[25.11/sqrt(20)] = 1/5.59 = 0.1789
p-value = P(z > 0.1789) = normalcdf(0.1789,100) = 0.4290
---
Since the p-value is greater than 5% a
screening program wouild not be warranted.
------
Cheers,
Stan H.